OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations eliminated today?

OpenStudy (anonymous):

430 720 1150 1300

OpenStudy (anonymous):

Assign variables: One-step equations: w Two-step equations: z Then: 3w - 2z = 1300 and: w + z = 1150. Solve for w.

OpenStudy (anonymous):

Solving is the part I need help on

OpenStudy (anonymous):

Okay, explanation follows:

OpenStudy (anonymous):

3w - 2z = 1300 w + z = 1150 Now what we want to do is solve for w, meaning we want z gone. How do we get rid of z? We put w in terms of z, meaning we put z on its own and w on the other side of the equation. Then: 3w - 2z = 1300 Subtract 3w from both sides: -2z = 1300 - 3w Divide by -2: z = 1.5w - 650 Next: w + z = 1150 Subtract w from both sides: z = 1150 - w Now we want to isolate w so that z is completely gone. Here's how: First we need to establish that our basis is correct: z = z We know z = z, so we can substitute: 1.5w - 650 = 1150 - w Great! z is gone. Now we can solve for w: 1.5w + w = 1150 + 650 2.5w = 1800 w = 1800/2.5 = 720.